The gravitational force between 2 objects is 54N. If the distance between them is tripled, what will the force be? 


1
Expert's answer
2021-08-11T09:54:36-0400

The law of Universal Gravitation says

"F=G\frac{m_1m_2}{r^2}\sim\frac{1}{r^2}"

Hence, if the distance between objects is tripled, the force will be

"F=\frac{54\:\rm N}{3^2}=6\:\rm N"